You are given a bottle containing 100 mL of a 50X TAE stock solution and are told to make 900 mL of 1X TAE solution. Describe what you would do.
Answer #1

C1V1= C2V2

C1= concentration of the stock solution

V1 - the volume of stock solution that is used to make the working solution

C2 - concentration of the working solution.

V2 - Final volume

For the given solution

C1 = 50X

V1= need to find out.

V2 = 900ml

C2 = 1X

50 * V1 = 1 = 900

CV1 = 900/50 = 18.

So we need to add 18 ml from 50X TAE and 882 ml water to make 900 ml of 1X TAE from 50X stock TAE